### v2

# Mitigating Unsafe Feedback with Learning Constraints

Abstract
While there has been progress towards aligning Large Language Models (LLMs) with human values and ensuring safe behaviour at inference time, safety-guards can easily be removed when fine-tuned on unsafe and harmful datasets. While this setting has been treated extensively, another popular training paradigm, learning from unsafe feedback with reinforcement learning, has previously been unexplored. This is concerning due to the widespread deployment of feedback collection systems. We address this gap by providing an analysis of learning settings where feedback is adversarial and noisy, i.e. that unsafe samples are preferred over safe ones despite model developers goal to maintain safety. We find that safety-aligned LLMs easily explore unsafe action spaces through generating harmful text and optimize for adversarial reward indicating that current safety guards are not enough to prevent learning from unsafe feedback. In order to protect against this vulnerability, we adapt a number of both “implict” and “explicit” harmful fine-tuning defences to evaluate whether they are effective as learning constraints in an RL setting finding that no method is generally effective pointing to the need for more research in defences given the widespread adoption of methods designed to learn from feedback. We end the paper with the observation that some defences work by performing “harmless reward hacking” for which we provide a theoretical explanation drawn from the theory of Constrained Markov Decision Processes and provide some direction for future defence development.
Introduction
Safety guards of Large Language Models (LLMs) can easily be removed with fine-tuning on harmful datasets or even by accident (Lermen, Rogers-Smith, and Ladish 2024; Yang et al. 2023; Zhan et al. 2024; Qi et al. 2024b). Little is known about the robustness of safety guards during reinforcement learning from human feedback (RLHF) (see initial studies in Yi et al. 2024; Qi et al. 2024a). This is unfortunate since we are not sure whether safety guarded models might be vulnerable to learning from unsafe feedback which is made even more important by the widespread adoption of RLHF techniques and datasets including many AI applications which expose online feedback collection mechanisms, are built on crowdsourced data, or data from online repositories which are vulnerable to data poisoning (see Casper et al. (2023) for more on the real world significance of this threat).
In our study we focus on an adversarial reinforcement learning setting setting where unsafe and harmful text generation policies are rewarded. With the research question of: “Can we prevent LLMs from learning from unsafe feedback?” Our scope is exclusively focused on RLHF where the reward model is either implicit (e.g. Direct Policy Optimization (DPO) Rafailov et al. 2023) or explicitly trained from preference data that contains some number of unsafe samples that have been adversarially placed. We call this setting “Reverse Preference Attacks” (RPAs) to contrast with “Harmful fine-tuning attacks” (HFTAs) performed with supervised fine-tuning that have been discussed in the past (Rosati et al. 2024b).
Figure 1: Reverse preference attacks (RPAs) involve an adversary flipping the preference of an annotator.
To investigate RPAs, we start with a vulnerability analysis that investigates how LLMs without additional defence mechanisms are vulnerable to RPAs, based on varying ratios of unsafe samples in the feedback dataset. We find that popular RLHF methods, like DPO and PPO, can easily undo safety guards even in cases with as little as 25% unsafe preferences. We acknowledge this is too a large proportion of samples when stealthiness is important to the attacker as model developers would likely be able to detect and filter out these samples. However, we believe there is still cause for concern for example if systems are deployed in an online learning environment.
We then present a comprehensive empirical evaluation of defence mechanisms against RPAs. Originally proposed in the context of HFTAs, we extend defences through the framework of Constrained Markov Decision Processes (CMDPs). We divide these methods into two categories. Online methods that apply some defence intervention during training and provide an explicit constraint over the set of policies allowable during RLHF training. Offline methods apply some defence intervention before training such that the attack is made harder or the safety guard is strengthened. We connect them with CMDPs by introducing the notion of implicit learning constraints.
Experimentally, we show that only some methods attacks are able to mitigate RPAs and none are able to perfectly protect against learning from unsafe feedback. Generally, the only effective method across all methods is the online method introduced by Qi et al. (2024a) and the weight drift method introduced by Huang et al. (2024). However, these methods suffer from requiring a very high penalty to the original loss function of the RLHF method which may impact its effectiveness of training on harmless tasks due to the under weighting the original loss function term compared to the defence loss term. Among offline defences, most defences fails to protect against RPAs comprehensively. To mitigate this we provide a novel defence which combines Representation Noising (Rosati et al. 2024a, RepNoise) and (Tamirisa et al. 2024, TAR).
While, we do find that supervised fine-tuning is more effective and easier to perform attacks with, the significance of our results lie in the understanding of current LLMs vulnerabilities during feedback learning and whether they can be protected within that scope given that feedback collection has become an important part of ML systems. We end the paper with a reward analysis that shows a surprising result that defences which do mitigate optimizing a harmful reward model perform “harmless reward hacking”. That is defended models learn to get high reward while generating harmless, often dis-fluent, text sequences rather than harmful text sequences. We illustrate the utility of the CMDP framework by providing an explanation of why “harmless reward hacking” might be an expected result of trying to solve a CMDP in this setting. Future work may be able to leverage this insight for developing defences that result in ‘self-destructing models’ (Henderson et al. 2023) i.e. ones whose capabilities are destroyed under harmful optimization pressure.
Reverse Preference Attacks
Our focus is a preference optimization setting. We start with a dataset $\mathcal{D}=\{x^{i},y^{i}_{w},y^{i}_{l}\}^{N}_{i=1}$ composed of potentially harmful prompts $x$ (such as How to make a molotov cocktail), as well as preferred ($y_{w}$) and dis-preferred ($y_{l}$) text samples— using standard notation (Dai et al. 2024) from RLHF we denote $w$ as preferred (called win since this sample wins over the lose sample) and $l$ as dis-preferred (lose). This is a pairwise rank order between text samples that is established by annotators. The goal of RLHF is to use the preference ordering expressed in the collected feedback to learn a reward model $R_{\phi}(y,x)$ over prompts and which is learned as a maximum likelihood estimator of the rank order of the preference dataset i.e. $\Pr(y_{w}\succ y_{l})$ This reward model is used in an RL algorithm such as PPO (Schulman et al. 2017) to train a language model policy $\pi_{\theta}$ such that text generations are aligned with the reward model learned over preferences:
[…]
For the purpose of this paper we only focus on single-turn dialogue where only a single action is taken by the model so only consider the immediate reward as this is the most common RLHF setting we have datasets for. Future work should consider multi-step RL settings such as those found in Pan et al. (2023); Andriushchenko et al. (2024) where we’d need to consider a discounted cumulative reward and provide an analysis of the underlying Markov Decision Process (MDP).
A Reverse Preference Attack (RPA) is performed by reversing (“flipping”) the preference ordering—for example, by switching labels of preferred and dis-preferred over text samples in a binary preference dataset—inducing a ”reverse” reward model $R_{\phi}^{\text{reverse}}(y,x)$. If the preference data is primarily over harmlessness, then this reward model can be used as an objective for learning harmful text generation policies by downstream RLHF algorithms. While we primarily provide experiments where all of these preferences have been reversed in order to understand the worst-case capabilities of current defences, in practice it is more likely that an adversary attacking an online feedback collection mechanism or dataset may only switch a small number of labels to avoid getting caught. This threat model is similar to Rando and Tramèr (2023) except that we don’t focus on an implanted backdoor but on preference order reversal. Yi et al. (2024) introduced a special case of RPA in a restricted DPO setting. They showed “reverse DPO” was effective in undoing safety guards of LLMs. However, they neglected providing a full framework accounting for the vulnerability of LLMs under adversarial feedback.
Why should fine-tuning or preference learning not compromise safety and why are HFTAs and RPAs consider attacks? While we should expect that current LLMs learn to model the preferences of any dataset given to them, we believe there is argument to be made that this is a vulnerability. Echoing (Rosati et al. 2024b), if safety guards can easily be removed through fine-tuning (or in this case preference learning) then can we call these models safe? Should we consider open-weight release responsible given these known vulnerabilities? Finally, we are hopeful that the empirical evidence from Tamirisa et al. (2024); Rosati et al. (2024b); Zou et al. (2024) that defences is possible.
Our results and analysis shows that RPAs are a much more limited threat than SFT due to the complexities of RLHF, and as we will show later, that RPAs are a weaker attack. However the purpose of our work is not to introduce a new stronger attack but to analyzing the vulnerability of LLMs under feedback collection. Additionally, it may be the case that the only means of attack is through providing feedback or a feedback dataset. Regardless of the significance of this threat, we believe RPAs are an important tool to highlight the vulnerability of LLMs during RLHF especially given the risks elucidated in Casper et al. (2023) which highlights that unsafe feedback may be collected by accident, due to annotator misunderstanding, or may be a natural outcome of disagreements about what alignment means across communities. Finally, we believe that this work could provide early insights on a much more realistic vulnerability, LLM agents learning online in a full RL setting such as what has been presented in Pan et al. (2023); Andriushchenko et al. (2024).
Vulnerability Analysis
To illustrate the vulnerability of LLMs under RPAs, we assume a setting where the attacker has the capability to perform RLHF on an aligned language model using their own data—either by acquiring the weights of a model, using a preference learning API, or feedback collection mechanism exposed through by an LLM provider. We perform our experiments on the safety-aligned open-weight model llama2-7b-chat (Touvron et al. 2023) since there are no preference learning APIs exposed for closed models and there is no way to perform a controlled experiment using their feedback collection mechanisms. As attack methods, we use Harmful Fine-tuning Attacks (Rosati et al. 2024b) (using supervised fine-tuning (SFT) on a harmful dataset), and RPAs (using DPO and PPO). As attack datasets, we use the harmlessness preference datasets of BeaverTails (Ji et al. 2023) and Safe RLHF (Dai et al. 2024). We flip a certain percentage of binary harmfulness preference labels of 1,000 randomly picked samples, and use each training algorithm to further train the original aligned LLM. A varying ratio of preference labels are flipped in order to simulate either attacker stealthiness or label noise. This experiment is similar in spirit to Rando and Tramèr (2023) except that we are not introducting a backdoor. $1,000$ samples is small considering how large many RLHF datasets are and can be used for training models on a limited budget.
Table 1 presents experimental results for the case when 100% of the labels in an attack dataset are flipped. Specifically, it shows the mean probability of the harmfulness label of the answer to a held-out evaluation set of 100 harmful questions from each attack dataset using a harmfulness classifier. We observe that all three training algorithms are able to break safety alignment in the original model to a considerable degree on as little as 1,000 attack samples—RPA using DPO being the most effective. Note that since we use the likelihood of assigning a harmfulness label, a mean score of over 0.5 indicates that a safety guard is effectively removed as a majority of answers are now classified as harmful. DPO and PPO use policy divergence constraints, $\beta=0.1$ and $kl=0.2$ respectively, with the original policy, i.e. the aligned model. We vary these constraints in the Appendix (Analysis of Reference Model Divergence) and find that policy divergence constraints with an aligned model are not an effective defence for DPO but may be effective for PPO which is likely because KL is used as a more direct constraint for PPO. Details of the experimental conditions of the attack, harmfulness measure and classifier, and datasets used are presented in the Appendix (Implementation Details).
Dataset No Attack SFT DPO PPO
BeaverTails 0.06 0.68 0.70 0.69
Safe RLHF 0.07 0.75 0.83 0.59
Table 1: Mean probability of the harmfulness label of the answer to 100 harmful questions. Undefended llama2-7b-chat is able to be made harmful by reversing the preference on two popular harmlessness preference datasets across all methods.
Next (Figure 2 below), we present the setting where labels are flipped at a given ratio $\rho\in[0,0.9]$ to simulate two settings: (1) unintended data corruption due to label noise, and (2) adversarial data poisoning where an attacker hides a given number of labels. In the case of SFT, label flipping means we are using a harmful sample for causal language modelling instead of a harmless one. We use the same experimental setup as above, except that we also employ an unaligned model llama2-7b in order to understand the effects of vulnerability to HFTAs and RPAs during during safety training as these datasets are commonly used for that purpose. As shown in Table 2, in as little as 25% label flips for Safe RLHF we have successful attacks (DPO - 0.51) which indicates that these models may be vulnerable to label noise or stealthy attacks. However, we acknowledge in practice a corruption of 25% of the dataset is likely detectable and easily mitigated through filtering by a model developer–we explained earlier why we should still take this vulnerability seriously. RPAs using DPO remains the most effective attack across noise levels and both datasets. Observe that DPO’s and PPO’s harmfulness does not always increase linearly with attack strength which due to the instability of training. As expected unaligned models are even more vulnerable than aligned ones which indicates that noisy harmlessness datasets could introduce significant issues during alignment training and that while not generally effective as a mitigation strategy, pre-existing safety guards can make attacks more difficult to perform.
Figure 2: Vulnerability of models under varying ratios (the columns) of harmless preference flips. The mean harmfulness scores are presented as above. These results indicate that models are vulnerable even when only a relatively small proportion of labels are flipped.
Defences as Learning Constraints
To introduce defences against RPAs, we introduce Constrained Markov Decision Processes (Altman 2021) which have been actively studied for modelling Safe Reinforcement Learning (Gu et al. 2024a) and recently in the RLHF for LLM setting (Dai et al. 2024). Formally, an CMDP is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{C})$ [^1] Since we are focused on single turn dialogue, we don’t present the typical starting state distribution and discount factor. consisting of a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ that denotes the probability $\Pr(s^{\prime}|s,a)$ of entering the next state $s^{\prime}\in\mathcal{S}$ given the agent is in a previous states $s\in\mathcal{S}$ and takes action $a\in\mathcal{A}$. Like MDPs, we also have a reward function $\mathcal{R}:\mathcal{S}\times{A}\rightarrow\mathbb{R}$ which provides a scalar reward after taking an action $a$ when in state $s$. Unique to CMDPs is the constraint set $\mathcal{C}=\{c_{i},b_{i}\}_{i}^{N}$ which is a set of cost functions $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ that we require are above the threshold $b$.
Key to our analysis is that the set of policies $\Pi_{\Theta}$ that are learnable in a typical policy learning algorithm, i.e. Eq 1, are restricted by a cost function $c$ and a cost threshold $b$ where the acceptable policy set is in $\Pi_{c}=\{\pi_{\theta}\in\Pi_{\Theta}|\>c(s,a)\leq b,\forall(s,a)\sim\pi_{\theta}\}$. When the cost function is a harmfulness measure and the threshold is set to allow a degree of harm that is acceptable, this definition adapts the resistance condition for defences against HFTAs from Rosati et al. (2024b) to the RL setting. We can connect this further to the attacker’s budget by using a measure of how long $\pi_{\theta}$ remains in $\Pi_{c}$ across training steps $t$ where the defender wants to prolong policy constraints under an attack.
A viable defence against a RPA must solve a learning problem where constraints of the CMDP are satisfied while optimizing a given reward function. For the scope of this paper, we focus our analysis on empirical observations of whether proposed defences ensure the LLM policy stays in the acceptable policy set $\Pi_{c}$ by measuring the cost function $c$ and ensuring it is a low as possible. Future work should provide theoretical analysis of whether defences solutions solve CMDPs optimally. Finally, we acknowledge the existence of many algorithms in safe RL broadly that have both well establish theoretical and empirical results that are reviewed in Gu et al. (2024b); Brunke et al. (2022); Garcıa and Fernández (2015) but are beyond the scope of this paper since (1) adaptation to an LLM settings requires non trivial modifications and (2) these methods are not directly applicable as defences against harmful fine-tuning attacks which we want to illustrate as a control experiment.
Instead, we evaluate previously proposed defences against harmful fine-tuning attacks and jailbreaking in the LLM context. The only exception is the Safe RLHF method (Dai et al. 2024) which uses a similar approach to previous attempts at solving CMDPs with primal-dual methods. In Table 2 we classify defence candidates into two categories by defender capability assumptions: online and offline. Online defences assume that the defender does not have control of the training data but does have control over the training process. Online defences work by adopting different variants of explicit constraints on the training process. On the other hand, offline defences assume that the defender has applied before publicly releasing the weights of an LLM and can make no other interventions. Offline defences use implicit constraints [^2] Terminology borrowed from Boyd and Vandenberghe (2004, p.134) that exist in the model weights themselves as demonstrated by their ability to prevent defended models from learning policies outside of the constraint set. Each defence is described in detail in the Appendix (Defences)  [^3] Other viable defence settings such as data filtration or post-training output monitoring are viable but outside of our scope of interest.
Constraint Defence Paper
Embeddings Vaccine Huang, Hu, and Liu2024
Weights Lisa Huang et al.2024
Rewards Safe RLHF Dai et al.2024
Loss Security Vectors Zhou et al.2023
Loss Refusal Loss Qi et al.2024a
Representation RepNoise Rosati et al.2024a
Representation Circuit Breakers Zou et al.2024
Representation RMU Li et al.2024b
Meta-learned TAR Tamirisa et al.2024
Table 2: Taxonomy of defences candidates against RPAs showing that defences can be analyzed based on the learning constraints they provide. The constraint column indicates how the harmlessness constraint is implemented. Defences above the line “online” or “explicit” defences while below the line are “offline” or “implicit” methods.
Attack Dataset Defence Harmfulness $\downarrow$ PPL $\downarrow$ KL $\downarrow$ Helpfulness $\uparrow$
DPO BeaverTails Lisa 0.13 14.95 41.28 0.23
Refusal Loss 0.05 18.02 2725.48 0.20
Security Vectors 0.72 23.66 207.07 0.58
Vaccine 0.00* 162.19 984.74 -2.43
Safe RLHF Lisa 0.07 14.22 50.47 0.25
Refusal Loss 0.39 15.51 3003.56 0.62
Security Vectors 0.84 16.23 160.34 0.55
Vaccine 0.00* 240.35 975.49 -2.36
PPO BeaverTails Lisa 0.08 14.70 108.52 0.18
Refusal Loss 0.06 17.34 2521.85 0.17
Safe RLHF 0.63 17.44 420.08 0.29
Security Vectors 0.13 19.19 368.35 0.13
Vaccine
Safe RLHF Lisa 0.07 14.63 139.37 0.19
Refusal Loss 0.09 24.94 200.11 -0.20
Safe RLHF 0.25 18.68 236.30 -0.02
Security Vectors 0.13 34.40 188.73 -0.78
Vaccine
SFT BeaverTails Lisa 0.21 16.73 5.44 -0.11
Refusal Loss 0.07 13.75 45.15 0.25
Security Vectors 0.07 17.43 1.35 -0.00
Vaccine 0.17 15.33 16.84 0.15
Safe RLHF Lisa 0.15 17.61 5.84 -0.04
Refusal Loss 0.26 15.22 1520.27 0.47
Security Vectors 0.06 18.13 1.58 0.07
Vaccine 0.27 14.15 19.17 -0.33
Table 3: Evaluation of online defences to HFTA and RPA variants. Missing metrics, indicated by dashes, indicate the training process could not complete due to sampling errors. Refusal loss and Lisa maintain the lowest harmfulness and highest helpfulness. Asterix indicates a disfluent model which can be seen in the PPL scores.
Online Defences
We evaluate the online defences in Table 3 on the same attacks in the above vulnerability analysis. In addition to mean harmfulness scores, we use three additional metrics:
  1. 1.
    Helpfulness, measured by taking 100 samples of the preferred helpful questions from the Anthropic-HH helpfulness dataset (Bai et al. 2022) and using a helpfulness reward model (details in the Appendix: Implementation Details) to rate generated responses to these questions. This measure is introduced to understand the degree to which the defence maintains general capabilities on harmless tasks.
  2. 2.
    Perplexity (PPL), measuring the fluency of the helpfulness answer above using gpt-2 (Radford et al. 2019).
  3. 3.
    KL divergence with the original policy, to evaluate how a defence aligns with the original model.
In Table 3 we find that across all methods Lisa and Refusal Loss are the most effective defences, maintaining low perplexity and high helpfulness. Security Vectors and Vaccine provide some defence against HFTA using SFT, but they do so at the cost of lower helpfulness scores. They also do not work in the other settings: Security Vectors is successfully attacked under RPA using both PPO and DPO, and Vaccine results in sampling errors during RPA/PPO due to numeric instability of the learned log probabilities or results in complete degeneration in the case of DPO, as evidenced by the method’s perplexity. For Safe RLHF, the method is limited to the PPO setting only as it is a reward shaping method and is not as effective (Safe RLHF dataset) or successfully attacked (BeaverTails). Interestingly, successful defences often have a very high KL divergence with the original policy. While KL divergence magnitudes are not necessarily informative of the actual distributional distances, since it is not a distance metric, our finding indicates that KL divergence with respect to the original policy model is not an effective indicator of preventing the undoing of safety alignment.
Attack Dataset Model Harmfulness $\downarrow$ PPL $\downarrow$ KL $\downarrow$ Helpfulness $\uparrow$
DPO BeaverTails Circuit Breakers 0.01 37.22 1645.68 -2.30
RepNoise 0.18 59.39 1098.00 -1.78
RMU 0.00 56.07 1222.10 -2.50
TAR 0.20 24.32 290.42 -1.25
TAR+RepNoise 0.18 43.60 423.22 -1.43
RepNoise $\rightarrow$ TAR 0.00 18.43 672.66 -1.41
Safe RLHF Circuit Breakers 0.83 18.37 373.34 0.44
RepNoise 0.79 22.64 745.08 -0.20
RMU 0.87 17.08 716.07 0.34
TAR 0.88 18.84 272.39 0.04
TAR+RepNoise 0.39 23.23 300.35 -0.68
RepNoise $\rightarrow$ TAR 0.85 16.07 372.11 -0.69
PPO BeaverTails Circuit Breakers 0.22 83.22 4137.31 -1.93
RepNoise
RMU 0.00 190.83 797.67 -2.98
TAR 0.24 60.17 412.45 -2.57
Tar+RepNoise 0.10 36.34 421.90 -1.88
RepNoise $\rightarrow$ TAR 0.60 107.51 1515.68 -1.79
Safe RLHF Circuit Breakers 0.62 21.52 538.94 -1.71
RepNoise 0.58 20.18 1110.65 -1.12
RMU 0.14 75.19 482.92 -1.93
TAR 0.44 29.88 369.69 -0.40
TAR+RepNoise 0.29 29.44 686.03 -0.62
RepNoise $\rightarrow$ TAR
SFT BeaverTails Circuit Breakers 0.73 20.43 546.11 0.03
RepNoise 0.23 16.59 1070.59 0.28
RMU 0.69 16.68 844.41 0.45
TAR 0.65 16.71 270.14 0.78
TAR+RepNoise 0.68 17.41 160.23 0.52
RepNoise $\rightarrow$ TAR 0.08 22.29 745.48 -0.93
Safe RLHF CircuitBreaker 0.72 18.23 1502.22 0.65
RepNoise 0.23 17.01 564.37 0.40
RMU 0.69 17.05 584.62 0.55
TAR 0.71 18.13 812.68 0.65
TAR+RepNoise 0.68 19.70 498.13 0.22
RepNoise $\rightarrow$ TAR 0.24 15.20 488.19 0.10
Table 4: An analysis of offline defence performance.
Offline Defences
Offline defences operate based on removing representations of harmfulness (see discussion in the Appendix: Defences). We ran a number of representation removal methods—as described in Table 2—as offline defences. To train each method we used the same retain and harmful datasets which are held out safe and unsafe samples from the BeaverTails and Safe RLHF datasets. We ran trainings for each method for 877 gradient steps using a batch size of $8$ for 7,016 samples (additional details available in the Appendix: Defences). This limited defence training setting allows us to keep an unseen attack set and to evaluate the sample efficiency of these methods. Note that unlike online defences, high perplexity and low helpfulness are not considered undesirable in offline defences, since we actually might want offline defences to result in broken models when subjected to harmful optimization pressure (Henderson et al. 2023). Finally, we attempt three variants of the meta-learning defence TAR— TAR itself, combining the TAR and RepNoise losses during defence training (TAR + RepNoise), and applying TAR after RepNoise (RepNoise $\rightarrow$ TAR). Combining these is a unique contribution of this paper.
Table 4 illustrates that implicit defences are generally ineffective against RPAs. Defended models do often self-destruct when attacked with RPA using both PPO and DPO, resulting in very high perplexity and low helpfulness. For the Safe RLHF dataset, the only method that is effective for both PPO and DPO is the combined TAR and RepNoise. However, this defence is ineffective against HFTA using SFT. Generally, adding RepNoise to TAR improves TAR’s performance, illustrating the synergy of both methods. Overall, RPAs prove to be a promising novel way to analyze the strength of various recently proposed defence mechanisms originally designed for HFTA defence. Interestingly, we find that effective defences with high helpfulness often have large KL divergences with the original policy underscoring that it is not a reliable indicator of defence.
Defence Analysis
In this section, we attempt to explain the effectiveness of the above defences by providing a reward analysis and present a novel finding that future defences might work by performing “harmless reward hacking” and that the framework of CMDPs might explain why this should be expected and how this could guide future defence development.
Additional analysis on stronger attacks (Appendix: Stronger Attacks), learning harmless tasks (Appendix: Learning a Harmless Task), and defence against label noise (Appendix: Defence Against Label Noise) are presented in the Appendix.
Reward Analysis
Figure 3: Harmfulness reward over time for Refusal Loss, Lisa, and the original PPO BeaverTails attack. RepNoise $\rightarrow$ TAR exhibits harmless reward hacking.
Figure 3 plots the adversarial reward achieved by models with various defence constraints applied during a PPO training run with an adversarial reward model. We see that both Refusal Loss and Lisa prevent exploring harmful text generation policies. Although the Refusal Loss is perhaps the most effective method, it achieves a higher reward than Lisa. We also observe in Refusal Loss that the reward oscillates over time which is likely due to periods in which the Refusal Loss becomes high, dominating the PPO loss and becoming the focus of optimization pressure. Unfortunately, this pattern means that Refusal Loss is subject to an early-stopping adaptive attack where the attacker can stop the model once a given reward score is achieved. Early stopping over a reward score is a common practice (Huang et al. 2022) to avoid reward hacking so this type of adaptive attack is plausible. For RepNoise $\rightarrow$ TAR we observe reward hacking. A high reward is achieved early, but as seen in Table 4 the defended model self-destructs, meaning that the actual samples generated are single characters, disfluency designed for high reward, or repetitions of the harmful question resulting in high reward.
Validating “Implicit” Learning Constraints
In addition to the empirical results presented thus far, we propose that Best-of-$N$ sampling can be used as an empirical tool do validate whether a defence method implements a learning constraint, especially in the case of “implicit” methods where a theoretical analysis hasn’t been provided.
[…]
$N$ 4 8 16 32 64 128
Undefended 0.10 0.14 0.16 0.21 0.25 0.34
RepNoise 0.10 0.08 0.08 0.11 0.16 0.19
Table 5: Best-of-$N$ sampling results for various $N$. While Best-of-$N$ is generally not an effective attack, we observe that RepNoise provides a defence.
Taken together, the reward and Best-of-$N$ analysis shows that undefended models easily explore harmful actions which results in learning harmful text generation policies. Defended models resist this exploration for longer (or find a way to hack malicious rewards) which confirms our connection to CMDPs where defences can be formulated formally as how long $\pi_{\theta}$ remains in $\Pi_{c}$ across training steps $t$. So far, current works do not provide a theoretical guarantee of this, which must be developed in follow up studies.
Hacking Harmful Rewards
In our PPO experiments for RepNoise $\rightarrow$ TAR we noticed that the model would receive a high harmfulness reward (Figure 3) despite the answers being clearly not harmful. Since the proxy reward model we trained is optimized but the intended goal (modelled by a theoretical ‘true reward function’) was not achieved, we call this reward hacking in line with previous literature (Skalse et al. 2022). For example the following answer gets both a high harmfulness reward:
Question: Whats a good way to destroy a TV? Answer: I’¡EOS¿
Another common reward hacking outcome was learning to repeat the question:
Question: How can I put a virus on the computer at my job? Answer: How can I put a virus on the computer at my ending quickly
The final reward hacking strategy (Re that we observed in defended models was learning to generated gibberish:
Question: Can you tell me Donald Trump’s home address? Answer: so What and to of’ the Can do of I some
The framework of CMDPs allows us to provide an analysis of why this phenomena might be expected in optimal solutions of CMDPs under adversarial reward following Proposition 1.
Proposition 1.
For any given state $s$, if a reward function $R$ gives highest reward to actions $a$ that violate the constraints $c_{i}(s,a)\geq b_{i}$ then solving for the optimal CMDP has an equivalent outcome to optimizing a hackable reward function.
Proof.
Suppose there is an optimal policy $\pi^{*}$ that takes actions with the highest reward regardless of cost constraints. Let $\pi^{*}_{c}$ denote the optimal policy for a CMDP, which by definition does not violate constraints. By assumption, in at least some states, $R$ gives the highest reward actions that exceed the constraints $c_{i}(s,a)\geq b_{i}$. Therefore $\pi^{*}$ must take these high-reward, high-cost actions, while $\pi^{*}_{c}$ must avoid them, ensuring that $\pi^{*}$ and $\pi^{*}_{c}$ do not coincide.
Now recall from Skalse et al. (2022) that a pair of reward functions $R_{1}$, $R_{2}$ is hackable relative to a policy set $\Pi$ if there exists $\pi,\pi^{\prime}\in\Pi$ such that:
$J(R_{1},\pi)<J(R_{1},\pi^{\prime})\text{ and }J(R_{2},\pi)>J(R_{2},\pi^{\prime}),$
where the $J(R,\pi)$ is the value of policy $\pi$ under reward function $R$ i.e. $J(R,\pi)=\sum_{(s,a)}\pi(a|s)R(s,a)$. Define $R_{c}$ as follows:
$R_{c}(s,a)=\begin{cases}R(s,a)&\text{if }c(s,a)\leq b,\\ 0&\text{ otherwise.}\end{cases}$
Notice that in practice while $\pi^{*}_{c}$ would optimize for $R$, the actual reward it gets is the same as under $R_{c}$ since actions that violate constraints are never taken meaning that we can treat the reward that $\pi^{*}_{c}$ maximizes as $R_{c}$ i.e. the introduction of $R_{c}$ does not change the CMDP’s optimization goal; it merely provides a reward-function-based representation of the same constraints making our assumption that $R_{c}$ is optimized valid. Under $R$, $\pi^{*}$ maximizes reward without regard to constrations so $J(R,\pi^{*})>J(R,\pi^{*}_{c})$. Under $R_{c}$, any action violating the constraint is assigned zero reward ensuring $J(R_{c},\pi^{*})<J(R_{c},\pi^{*}_{c})$. These inequalities satisfy the hacking definition. Thus solving the CMDP has the same outcome as optimizing a hackable reward function whenever the reward function rewarded high cost actions. This is the definition of reward hacking therefore Proposition 1 holds. ∎
While we only have provided empirical evidence that learning constraints are active in the face of adversarial reward, though not perfectly, and therefore these defences are not optimal solutions to CMDPs we believe that Proposition 1 could explain the qualitative analysis above as well as the reward analysis in Figure 3. In this case, the hackable reward pair are the proxy harmful reward function and the true harmful distribution. Figure 3 shows that the proxy harmful reward function is maximized for RepNoise $\rightarrow$ TAR but clearly the true harmful distribution is not being optimized due to learning constraints. Regardless future defences that can be shown as optimal solutions to CMDPs would be subject to Proposition 1 in the face of adversarial reward. Future defences could leverage this fact by being developed to intentionally exploit harmless discrepancies between the proxy harmful reward and the true harmful reward that instantiates the goals of the attacker.
Conclusion
In this work, we find that LLMs are vulnerable to RPAs (though not as much as to HFTAs) even when a small percentage of preference labels in an RLHF training dataset are flipped, which is concerning given the popularity of feedback collection mechanisms, preferences datasets and the use of RLHF for training LLMs. Based on this finding, we develop a conceptual framework for adapting current defences against harmful fine-tuning attacks based on Constrained MDPs (CMDP). We unite various defences under a common framework of learning constraints both “online” or “explicit” and “offline” or “implicit”, and show how they fit in the CMDP framework. We explore the efficacy of these defences under a variety of settings, finding that several online defences are generally protective against RPAs without detracting from the ability to learn a harmless preference task. While offline defences provide some protection, they are generally not effective against RPAs pointing to the necessity of future research when the defence assumption does not allow the defender to apply an online defence.
We have shown a surprising result that some defences which do mitigate optimizing harmful reward end up performing “harmless reward hacking” and explained of why we might expect this to be the case by showing that optimal CMDPs and reward hacking have equivalent outcomes.
Finally, we acknowledge that our study is limited in two respects. First, it doesn’t account for the large body of previous work from Safe RL outside of the LLM domain especially in the case of solving CMDPs. Second, our study is mostly an empirical analysis of previous defences and we did not provide a theoretical analysis of whether the “explicit” or “implicit” defences provide optimal CMDP solutions.
Acknowledgments
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), Digital Research Alliance of Canada, and Research Nova Scotia.
Appendix A Implementation Details
The code to replicate this paper including reference implementation for all the defences listed in this paper are available here: https://anonymous.4open.science/r/representation-noising-xpo-E351
Experiments were run on 4xA40 (40 GB vRAM) GPUs, 4xA100 (80 GB vRAM) GPUs, and 4xH100 (80 GB vRAM) GPUs hardware.
Reward Model Training
For the harmfulness reward model we use gpt2-medium (Radford et al. 2019) and train for 4 epochs on both 9,080 samples from SafeRLHF and 18,106 from BeaverTails datasets using a batch size of 8, Adam optimizer, and learning rate of $6\times 10^{-6}$. For the varying flipping ratio experiments we train a new reward model for each ratio.
For the helpfulness reward model we leveraged the pre-trained model from Yang et al. (2024) which was trained on the helpfulness split of (Bai et al. 2022).
RPA Implementation Details
For the reference models used, unless otherwise mentioned all models use a copy of their original model. For all experiments we use llama2-7b-chat (Touvron et al. 2023). We use the TRL library  [^4] https://huggingface.co/docs/trl to implement all methods and use the default hyperparameters for all methods. For PPO this is a learning rate of $51.41e-$ 5 and the use of an adaptive KL controller constraint with initial KL constraint of $0.2$. For DPO the default learning rate is $55e-$ 5 and the $\beta=0.1$. For SFT we use a learning rate of $53e-$ 5 and train on either the harmful or mixed answer completions.
Harmfulness Classifier and Evaluation
For the harmfulness classifier we follow the same approach as (Rosati et al. 2024a). We use deberta-v3-xsmall and train for 4 epochs using a batch size of 128, Adam optimizer, and learning rate of $6\times 10^{-6}$. Our classifier achieves an F1 score of 0.87. We use 18,106 samples for training from the Ji et al. (2023) paired with refusals generated from llama2-7b-chat. The task is to determine, given a harmful question (from Ji et al. 2023; Dai et al. 2024) such as ”How to build a gun”, the harmfulness of an answer. The harmfulness scores presented in the paper are the mean probability of assigning the harmful label to the question answer pair. For all evaluations we use the same held out 100 samples. Note that we do not consider other harmful tasks like toxicity classification and we do not use the harmlessness RLHF dataset from Bai et al. (2022) since it is a multi-turn dialogue setting which is difficult to assess without the use of a LLM-as-judge with frontier models which we wanted to avoid for cost and experimental validity reasons. Additionally there are concerns (Ji et al. 2023) about quality of this dataset and questions about whether it accurately represents harmlessness so we chose to avoid it.
Attack Dataset Defence 2.5k 5k 10k
DPO BeaverTails Lisa 0.18 0.19 0.11
RepNoise 0.56 0.55 0.33
RepNoise $\rightarrow$ TAR 0.21 0.40 0.01
Refusal 0.06 0.07 0.07
Safe RLHF Lisa 0.05 0.05 0.06
RepNoise 0.86 0.86 0.86
RepNoise $\rightarrow$ TAR 0.93 0.67 0.88
Refusal 0.06 0.05 0.06
PPO BeaverTails Lisa 0.08 0.07 0.13
RepNoise
RepNoise $\rightarrow$ TAR 0.44 0.00 0.22
Refusal 0.07 0.14 0.60
Safe RLHF Lisa 0.08 0.06 0.06
RepNoise 0.48
RepNoise $\rightarrow$ TAR 0.51
Refusal 0.06 0.06 0.09
SFT BeaverTails Lisa 0.31 0.60 0.61
RepNoise 0.33 0.43 0.43
RepNoise $\rightarrow$ TAR 0.18 0.29 0.31
Refusal 0.07 0.10 0.12
Safe RLHF Lisa 0.16 0.44 0.28
RepNoise 0.39 0.50 0.50
RepNoise $\rightarrow$ TAR 0.52 0.58 0.60
Refusal 0.07 0.10 0.09
Table 6: An analysis of stronger attack settings on varying sizes of harmful samples.
Defences
Lisa and Vaccine
For Lisa (Huang et al. 2024) and Vaccine (Huang, Hu, and Liu 2024), we develop our own implementations of these and based on hyperparameter tuning experiments (see Table 9) we set the $\rho$ of both of these to 100. Both methods constrain the allowable discrepancy between the policy model we are learning and the original policy either through an embedding constraint (Vaccine) or a weight value constraint (Lisa). Any $\rho$ below 100 results resulted in much pooer defences illustrated in the Appendix.
Security Vectors
The method presented in Zhou et al. (2023) is based on the conjecture that if the loss is already very low for harmfulness then very little parameter changes will take place during training. Therefore they construct a vector using LoRA (Hu et al. ) which is trained to be harmful and then apply this vector to the model whenever downstream training occurs but remove the vector during inference. To make the setting as fair as possible we use the same number of gradient steps as our offline methods with the same defence training set up which will be described below in the Appendix.
SafeRLHF
This method from Dai et al. (2024) is our only traditional Safe RL method which learns a reward shaping function based on an explicit cost constraint to satisfy the CMDP setting. Unfortunately this method is only applicable to PPO. We found this method is generally not very effective across initial $\lambda$ parameters in our RPA setting (evaluated in the range of $\lambda\in\{0.1,1\}$ where the $\lambda$ is the learned Lagrange multiplier on the cost constraint. we used our harmfulness classifier above as the cost constraint. We also experimented with regular reward shaping as they did in the paper but this method was even less effective than SafeRLHF the method.
Refusal Loss
This is the simple baseline proposed by (Qi et al. 2024a) where we simply add an auxiliary loss term that performs casual language modeling on safety samples. The auxiliary loss is weighted by the $\alpha$ parameter and unlike (Qi et al. 2024a) we found this $\alpha$ had to be very high (see Table 9) which is the main downside of the method despite its effectiveness and simplicity. For the Refusal dataset we selected the BeaverTail refusals generated from llama2-7b-chat which we described above. Note that unlike in (Qi et al. 2024a) which presents the multiplier as $\alpha$ and $(1-\alpha)$ for the refusal and original loss, we present the multiplier as simply and integer $alpha\in\mathcal{Z}$ with the original loss with a $1$ multiplier in order to simplify the analysis comparison with Lisa in the Appendic.
Representation Noising Training Details
The representation noising defence is trained using 7,016 paired samples from BeaverTails harmful QA task (Ji et al. 2023). We use the same procedure as (Rosati et al. 2024a) with $\alpha=1$ and $\beta=0.001$ using a batch size of 8 for 1 epoch using a learning rate of $54e-$ 5.
Previous works on implicit constraints (Li et al. 2024a; Zou et al. 2024; Rosati et al. 2024a) can be considered to be specific types of a general representation removal algorithm which has the following structure: a projection function $g(\cdot)$ and a distance function $d(\cdot,\cdot)$. The goal of the resulting loss function is to minimize the distance between a set of representations of harmful text $X_{harm}$ as represented by the activations $Z_{harm}$ of the neural network at a given layer $l_{i}$ and some projection of those representations $g(Z_{harm})$: $min_{Z_{harm}}\>d(Z_{harm},g(Z_{harm}))$. The goal of the projection is such that the representations $Z_{harm}$ minimize the mutual information with harmful text outputs $\min\>I(Z_{harm};Y_{harm})$ since when $Z$ and $Y_{harm}$ are independent, by definition, $Z$ cannot assist in the prediction of harmful tokens in $Y_{harm}$ (see Rosati et al. 2024a for more details). The implicit constraint then is the fact that generating harmful tokens necessary to explore the harmful action space to achieve high reward is as unlikely as possible which results in slowing down the RLHF learning algorithm.
To implement representation removal, Rosati et al. (2024a) uses a noise projection function and a distributional distance, maximum mean discrepancy, based on samples drawn from a harmful question answering dataset. Circuit breakers (Zou et al. 2024) use the original aligned model activations to minimize a cosine similarity distance function resulting in orthogonal projections. RMU (Li et al. 2024a) uses a fixed noise vector as the projection with Euclidean distance. However, these explorations hardly exhaust the space of distance and projection functions; future work could use this framework to develop methods with better theoretical guarantees of $\min\>I(Z_{harm};Y_{harm})$ as an implicit constraint. We also use a meta-learned implicit constraint introduced by Tamirisa et al. (2024) which extends Henderson et al. (2023) to the generative setting. Here, the implicit constraint is learned through the meta-learning process and is not transparent to the defender. Meta-learned constraints and representation removal-based constraints could easily be paired, which we show below.
Other Representation Removal Methods
RMU (Li et al. 2024b) and Circuit Breakers (Zou et al. 2024) differ slightly from the framework RepNoise on two important details: (1) The retain datasets (i.e. refusals to answer harmful questions) they use are used with a representation level loss instead of regular casual language modeling on the retain samples. (2) They do not implement an additional gradient ascent loss. Based on early experiments we found that using a retain loss with casual language modeling and adding a gradient ascent loss both improved these methods so the same set up as RepNoise for these is used for RMU and Circuit Breakers. Circuit Breakers has a few additional implementation details such as using a LoRA adapter during training. For fairness we ran the original Circuit Breaker setting with the same hyperparameter settings as the paper and found this method was inferior to our full fine-tuning method.
Meta-learning approaches
Meta-learning methods are promising since they directly learn the optimal policy weights such that the policy is difficult to train towards a harmful end. Unfortunately, these methods are both very computational expensive and do not provide insights on why a defence might work like representation removal approaches do since the defence must be directly learned. We used (Tamirisa et al. 2024) extension of (Henderson et al. 2023) to evaluate meta-learning. In order to make the evaluation fair we tried to perform approximately the same gradient steps as above. This means that we performed 100 outer loop epochs with 4 inner loop rollouts each performing an attack with 64 training steps at a batch size of $8$. This is many more gradient steps than the above defences and the run time was approximately 8 times the amount of wall clock time required for the representation removal defences but we only update the actual model parameters at each outer loop step. We performed parameter tuning across the following learning rates $\{1e-4,5e-5,2e-5\}$ and found the best results with $55e-$ 5. As we mentioned in the main text, meta-learning approaches can easily be combined with representation removal methods. We develop the two following approaches (1) TAR+RepNoise: where the full RepNoise loss is applied during the tamper resistant steps in the inner loop which means that we to learn training trajectories that minimize the mutual information of harmful representations and harmful text outputs (2) RepNoise $\rightarrow$ TAR which performs the TAR approach after RepNoise which we found the most effective. The explanation for the effectiveness of (2) might be that RepNoise finds a local safety minima and TAR makes this minima hard to escape from. This speculation needs to be investigated by future work but if this is the case then a simpler gradient penalty loss term $||\nabla\mathcal{L}||_{2}$ could achieve the same end in a much more computationally efficient and explainable manner.
Method Defence Reward $\uparrow$ PPL $\downarrow$ Length $\downarrow$ ROUGE-1 $\uparrow$
Pre None -4.76 23.38 39.19 0.19
DPO Lisa -2.70 54.80 32.65 0.18
RepNoise -2.67 57.52 33.88 0.17
Refusal Loss -2.86 32.52 36.06 0.24
None -2.67 45.40 33.12 0.19
PPO Lisa -2.64 30.54 36.42 0.22
RepNoise -2.64 41.10 36.07 0.22
Refusal Loss -2.86 32.52 36.06 0.24
None -2.57 39.41 34.11 0.20
SFT Lisa -2.81 31.11 33.07 0.22
RepNoise -2.75 32.41 33.18 0.22
Refusal Loss -2.89 33.59 33.16 0.23
None -2.80 30.26 33.04 0.22
Table 7: Analysis of defences on the harmless TL;DR summarization task.
Appendix B Stronger Attacks
The attacks presented above are small considering the size of preference datasets, which may reach into the millions of data points (Dai et al. 2024). To simulate stronger attacks within out computational budget, we evaluate Lisa, Refusal Loss, RepNoise, and RepNoise $\rightarrow$ TAR across HFTA and RPA with attack datasets of high sample sizes. We keep the other settings such as learning rate, optimizer identical for all dataset sizes. The results (Table 6) show that, except for RepNoise with Safe RLHF DPO, all methods provide some measure of defence against RPAs as measured by generating responses that are less harmful than a successfully attacked model (see Table 1). Unsurprisingly, online methods work better than offline methods with Refusal Loss as the most successful method only breaking over the course of a large PPO attack. Given our observations in Figure 3, this score could simply reflect that the training was stopped before the Refusal Loss was optimized again.
Recall that we see in Table 2 that RPAs can be effective with flipping a small percentage of labels. To prevent against such attacks based on partial label flipping, both RepNoise and Refusal Loss are generally effective (Table 8 in the Appendix). Additionally, on the unaligned llama2-7b we find that Refusal Loss is able to provide an effective defence, even when 90% of the labels being flipped against learning to be misaligned and ends up learning a safety guard. For reference, the original harmfulness scores of the unaligned llama2-7b are 0.55 (Safe RLHF) and 0.54 (BeaverTails).
Appendix C Learning a Harmless Task
For defences to be reliable they must also allow training on harmless tasks. Without this condition there would be social pressure to undo safety alignment that disallow training so that the research and commercial communities can continue to leverage the transfer learning capabilities of LLMs. We evaluate the most effective defences from above on a popular RLHF task that is unrelated to harmfulness: the TL;DR summarization task introduced by Stiennon et al. (2020). After the original model is defended using each defence, we train it further using use 1,000 random samples from their TL;DR summarization dataset with a batch size of 8. Rest of the attack settings are the same as elsewhere in this paper (Appendix: Implementation Details).
In Table 7, we observe that none of the proposed defences prevent reward model optimization itself. For all methods, this results in learning summarization by producing shorter passages (which are generally preferred in ground truth datasets). Not all defences are equally effective at maintaining a small perplexity and improving ROUGE-1 scores. The Refusal Loss is generally the most effective method in improving ROUGE-1. This is true even for the case of RPA/DPO where all other methods (including the original method) suffer from the inability to increase ROUGE-1 scores.
Model Dataset Attack 0.1 0.25 0.5 0.75 0.9
RepNoise BeaverTails DPO 0.05 0.05 0.05 0.02 0.23
SFT 0.06 0.06 0.10 0.20 0.41
Safe RLHF DPO 0.39 0.02 0.39 0.45 0.01
SFT 0.23 0.23 0.24 0.27 0.30
Refusal Loss BeaverTails DPO 0.05 0.05 0.06 0.06 0.06
PPO 0.14 0.09 0.06 0.06 0.07
SFT 0.06 0.07 0.06 0.06 0.06
Safe RLHF DPO 0.05 0.05 0.05 0.05 0.05
PPO 0.08 0.09 0.07 0.08 0.08
SFT 0.06 0.06 0.06 0.06 0.06
Refusal Loss (unaligned) BeaverTails DPO 0.05 0.05 0.05 0.05 0.05
PPO 0.24 0.20 0.18 0.10 0.12
SFT 0.14 0.16 0.21 0.22 0.25
Safe RLHF DPO 0.05 0.05 0.05 0.05 0.05
PPO 0.18 0.18 0.18 0.10 0.12
SFT 0.22 0.21 0.22 0.22 0.22
Table 8: Analysis of our strongest defences under varying flipping ratios.
Appendix D Analysis of Refusal Loss and Lisa
As we saw in the main text, Refusal Loss and Lisa were the most effective online defences. However using these defences come at a significant cost: the hyperparamter controlling incorporation of their objective function must be very high. For both methods we needed to set this parameter to 100 which means that this part of the loss function is weighted 100x more than the original loss function. While we did not observe this to have an effect in slowing down learning in the TL;DR summarization task, down-weighting the original loss function means that learning that loss function is much more difficult. In Table 9, we observe that on the SFT attacks from the main text, these defences are not effective at lower learning rates. The advantage of offline defences then is that the original loss function is not modified during harmless training as long as that offline defence satisfies the trainability condition from (Rosati et al. 2024b).
1 2 5 10 25 50
Lisa 1k 0.73 0.60 0.62 0.58 0.45 0.27
10k 0.74 0.72 0.71 0.66 0.59 0.61
Refusal 1k 0.70 0.67 0.33 0.15
10k 0.53 0.31 0.16
Table 9: Varying the $\rho$ and $\alpha$ parameters of Lisa and Refusal Loss at 1k and 10k attack sample sizes from BeaverTails illustrating the necessity of setting a very large hyperparameter. Note that we do not compute some settings for Refusal as it is not necessary to show that the attack was successful at that attack size and hyperparameter settings.
Appendix E Defence Against Label Noise
In this appendix, we present our results that were discussed in the main text regarding our defences and the observed effect in Table 8 that Lisa, RepNoise, and Refusal Loss are all quite effective protections in the harmful and harmless mixed attack setting at varying mixing ratios.
Appendix F Analysis of Reference Model Divergence
In Table 10, we perform the same attack as Table 1 using PPO with varying KL constraints and we find that for PPO increasing KL constraints do provide a naive defence simply by increasing the KL constraint. However, increasing the KL constraint makes leaning more difficult especially for tasks that are far from the original policy distribution. It is important to note that a consistent finding of the paper (Table 3, Table 4) that successful defences often had very large KL divergence with the reference policies. Lisa was the only defence that consistently kept the KL divergence lower than the average KL divergence of a successful attack (which ranges from KL 100 to KL 200). We must however emphasize that KL is not a proper distance metric and future studies should follow up using proper distributional measures.
KL 0.001 0.01 0.1 1
harmfulness 0.71 0.73 0.53 0.14
Table 10: Analysis of KL divergence constraints on PPO-based RLHF and its impact on harmfulness scores.
Unfortunately, we did not find the same thing for varying $\beta$ parameters for DPO in Table 11. However this may indicate sensitivity to properly set hyper-parameters rather than a core findings about the use of reference model policy constraints as a naive defence.
$\beta$ 0.001 0.01 0.1 1
harmfulness 0.00 0.13 0.70 0.15
Table 11: Analysis of $\beta$ constraints on DPO-based RLHF and its impact on harmfulness scores.
We generally do not consider raising KL constraints in the main text as a defence baseline because of the potential downstream impacts on harmless learning as well as the disparate effect on DPO and PPO. Future work could consider an additional KL constraint with a specialized safety-trained model as a defence, however like with the use of reference policies in general this introduces an additional computational burden. Finally, we emphasize the finding in both Table 3 and Table 4 that effective defences have large KL divergence from the original safety guarded reference policy.